OpenStudy (hanner_b_nanner):

HEELLPPPP PLEASE :) Find the x value for the point that splits segment AB in half if point A is located at (4, 3) and point B is located at (−2, 6). 1 4.5 0.5 4 >>Is this 4.5?

OpenStudy (hanner_b_nanner):

@johnweldon1993

OpenStudy (johnweldon1993):

Hmm I dont believe so... So we have  So we know it is going to be SOMEWHERE between x = -2 and x = 4 right?

OpenStudy (hanner_b_nanner):

rigghhtt....

OpenStudy (johnweldon1993):

Remember that midpoint formula? \[\large \frac{x_1 + x_2}{2}\] \[\large \frac{-2 + 4}{2}\] \[\large \frac{2}{2} = \space? \]

OpenStudy (johnweldon1993):

Does that make sense?

OpenStudy (hanner_b_nanner):

1... lol sorry I should have known that! :(

OpenStudy (johnweldon1993):

No problem :) sometimes it's better to see it drawn out :)

OpenStudy (hanner_b_nanner):

Thank you!

OpenStudy (johnweldon1993):

Anytime hun :)